# latexOverwriting file

### v1

@@ 3 Particle Filter
During the last quarter of the last century, the particle filter became a widely adopted methodology for solving the filtering problem in moderate dimensions $d$; the methodology is overviewed in [6].
@@
$\mathsf{S}^{J}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}=\frac{1}{J}\sum_{j=1}^{J}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\delta}_{v_{j}}\,,\qquad v_{j}\overset{\rm{i.i.d.}}{\sim}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}\,.$
@@
Since $\mathsf{S}^{J}\approx I$, we may introduce the approximation ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{n}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ evolving according to the map
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{n+1}$ $=\mathsf{L}(\mathsf{S}^{J}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{n};{y}^{\dagger}_{n+1}),\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{0}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{n}=\sum_{j=1}^{J}w_{n}^{(j)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\delta}_{\widehat{v}_{n}^{(j)}}$
@@
$\widehat{v}_{n+1}^{(j)}$ $=\Psi\bigl{(}v_{n}^{(j)}\bigr{)}+\xi_{n}^{(j)},\qquad v_{n}^{(j)}\overset{\rm i.i.d.}{\sim}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{n},$
@@
Systemization of the analysis of this method may be found in the work of Del Moral [4], with the significant extension to analysis over long time-intervals in [5]. We describe a prototypical theoretical result, based on the formulation underpinning Rebschini and Van Handel [17]; for precise statement of conditions under which the theorem holds see Law et al [14, Theorem 4.5]. The theorem deploys a metric $d(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ on random probability measures. This is defined as follows:
$d({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\nu}})^{2}=\sup_{|f|\leqslant 1}\mathbb{E}\bigl{|}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}[f]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\nu}}[f]\bigr{|}^{2},$
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}[f]=\int f(u)\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}(\mathrm{d}u)$ and similarly for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\nu}};$ the expectation denoted $\mathbb{E}$ is over the randomness used to define random probability measures. The metric reduces to TV distance if applied to deterministic probability measures.
@@
It holds that
$\sup_{0\leqslant n\leqslant N}d({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{\rm PF}_{n})$ $\leqslant\frac{C(N)}{\sqrt{J}}.$
@@
  • Consistency: $d(\mathsf{S}^{J}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}})\leqslant\frac{1}{\sqrt{J}}.$ This shows that $\mathsf{S}^{J}$ approximates the identity.
@@
In practice, especially in high dimensions $d$, the method often suffers from weight collapse [1, 20]. This refers to the phenomenon where one of the weights $\{w_{n}^{(j)}\}_{j=1}^{J}$ is close to $1$ and the others are close to $0$ When this happens the method is of little value because the effective number of particles approximating ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ is $1$ This leads us to consider the ensemble Kalman filter which, by design, has equal weights.
@@ 4 Ensemble Kalman Filter
The ensemble Kalman filter, introduced by Evensen in [8], is overviewed in the texts [7, 9]. We will state the basic particle form of the algorithm, from [8]. We will then show how it may be derived from a mean-field perspective highlighted in [2]. Finally we describe analysis of this mean-field ensemble Kalman filter, and its relationship to the true filter [3].
To describe the particle form of the method we first introduce key notation relating to covariances. Specifically we write covariance under ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})$ as:
$\mathcal{\rm cov}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})=\begin{pmatrix}\mathcal{C}^{vv}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})&\mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})\\ \mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})^{\top}&\mathcal{C}^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}})\end{pmatrix}.$
We use similar notation (${\rm mean}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}}$) for the mean under ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}.$ The ensemble Kalman filter from [8] then has the form, for $n\in\mathbb{Z}^{+}$,
@@
$v_{n+1}^{(j)}$ $=\widehat{v}_{n+1}^{(j)}+\mathcal{C}^{vy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1}\bigr{)}\mathcal{C}^{yy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1}\bigr{)}^{-1}\bigl{(}{y}^{\dagger}_{n+1}-\widehat{y}_{n+1}^{(j)}\bigr{)},$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1}$ $=\frac{1}{J}\sum_{j=1}^{J}\delta_{\bigl{(}\widehat{v}_{n+1}^{(j)},\widehat{y}_{n+1}^{(j)}\bigr{)}}.$
@@
Note that the components of ${\rm cov}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ involve empirical learning of the covariances $\Sigma$ and $\Gamma$ of the $\{\xi_{n}^{(j)}\}_{j=1}^{J}$ and $\{\eta_{n+1}^{(j)}\}_{j=1}^{J}$ respectively. It is possible, and indeed often desirable, to directly input the matrices $\Sigma$ and $\Gamma$, only using empirical estimation for the covariance based on the $\{\Psi(v_{n}^{(j)})\}_{j=1}^{J}$ and $\{h(\widehat{v}_{n+1}^{(j)})\}_{j=1}^{J}.$ In particular to use
$\mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ $\approx C^{vh,J},$ (4.1a)
$\mathcal{C}^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ $\approx C^{hh,J}+\Gamma$ (4.1b)
where $C^{vh,J}$ is the empirical cross-correlation between the $\{\Psi(v_{n}^{(j)})\}_{j=1}^{J}$ and the $\{h(\widehat{v}_{n+1}^{(j)})\}_{j=1}^{J}$, whilst $C^{hh,J}$ is the empirical correlation of the $\{h(\widehat{v}_{n+1}^{(j)})\}_{j=1}^{J}$. The actual definition of $C^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ instead uses an empirical approximation $\Gamma_{n+1}$ of $\Gamma$ formed from $\{\eta_{n+1}^{(j)}\}_{j=1}^{J}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{{\rm EK},J}_{n}=\frac{1}{J}\sum_{j=1}^{J}\delta_{v_{n}^{(j)}},$
noting that for $n=0$ we choose $v_{0}^{(j)}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{0}$ i.i.d. . Practitioners like this methodology because it assigns equal weights to the particles and cannot suffer from the weight collapse arising in the particle filter. Furthermore, in the setting where $\Psi,h$ are both linear, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ is Gaussian, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}^{{\rm EK},J}_{n}$ converges to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ as $J\to\infty$, at the Monte Carlo rate [15, 16]. In this linear setting, with Gaussian noise and initial condition, the problem can be solved explicitly by the Kalman filter [11]; this is a nonautonomous and nonlinear dynamical system for the mean and covariance of the (in this setting) Gaussian ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}.$ However forming and propagating $d\times d$ covariances, when $d\gg 1$, is impractical; ensemble methods, on the other hand, operate by computing a $J\times d$ low-rank approximation of the covariance and can be used when $d\gg 1.$ Thus particle methods are of some practical value even in this linear Gaussian setting, avoiding the need to work with large covariances when $d\gg 1.$

### v3

@@ 3 Particle Filter
During the last quarter of the last century, the particle filter became a widely adopted methodology for solving the filtering problem in moderate dimensions $d$; the methodology is overviewed in [doucet2001sequential].
@@
$\mathsf{S}^{J}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}=\frac{1}{J}\sum_{j=1}^{J}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta}_{v_{j}}\,,\qquad v_{j}\overset{\rm{i.i.d.}}{\sim}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}\,.$
@@
Since $\mathsf{S}^{J}\approx I$, we may introduce the approximation ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{n}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ evolving according to the map
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{n+1}$ $=\mathsf{L}(\mathsf{S}^{J}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{n};{y}^{\dagger}_{n+1}),\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{0}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{n}=\sum_{j=1}^{J}w_{n}^{(j)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\delta}_{\widehat{v}_{n}^{(j)}}$
@@
$\widehat{v}_{n+1}^{(j)}$ $=\Psi\bigl{(}v_{n}^{(j)}\bigr{)}+\xi_{n}^{(j)},\qquad v_{n}^{(j)}\overset{\rm i.i.d.}{\sim}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{n},$
@@
Systemization of the analysis of this method may be found in the work of Del Moral [del1997nonlinear], with the significant extension to analysis over long time-intervals in [del2001stability] and to continuous time in [cricsan1999interacting]. We describe a prototypical theoretical result, based on the formulation underpinning Rebschini and Van Handel [MR3375889]; for precise statement of conditions under which the theorem holds see Law et al [MR3363508, Theorem 4.5]. The theorem deploys a metric $d(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$ on random probability measures. This is defined as follows:
$d({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\nu}})^{2}=\sup_{|f|\leqslant 1}\mathbb{E}\bigl{|}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}[f]-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\nu}}[f]\bigr{|}^{2},$
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}[f]=\int f(u)\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}(\mathrm{d}u)$ and similarly for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\nu}};$ the expectation denoted $\mathbb{E}$ is over the randomness used to define random probability measures. The metric reduces to TV distance if applied to deterministic probability measures.
@@
[MR3363508, Theorem 4.5] It holds that
$\sup_{0\leqslant n\leqslant N}d({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{\rm PF}_{n})$ $\leqslant\frac{C(N)}{\sqrt{J}}.$
@@
  • Consistency: $d(\mathsf{S}^{J}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}})\leqslant\frac{1}{\sqrt{J}}.$ This shows that $\mathsf{S}^{J}$ approximates the identity.
@@
In practice, especially in high dimensions $d$, the method often suffers from weight collapse [MR2459233, snyder2008obstacles]. This refers to the phenomenon where one of the weights $\{w_{n}^{(j)}\}_{j=1}^{J}$ is close to $1$ and the others are therefore necessarily close to $0$ When this happens the method is of little value because the effective number of particles approximating ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ is $1$ This leads us to consider the ensemble Kalman filter which, by design, has equal weights.
@@ 4 Ensemble Kalman Filter
The ensemble Kalman filter, introduced by Evensen in [evensen1994sequential], is overviewed in the texts [evensen2009data, evensen2022data]. We will state the basic particle form of the algorithm, from [evensen1994sequential]. We will then show how it may be derived from a mean-field perspective highlighted in [2022arXiv220911371C]. Finally we describe analysis of this mean-field ensemble Kalman filter, and its relationship to the true filter [carrillo2022ensemble].
To describe the particle form of the method we first introduce key notation relating to covariances. Specifically we write covariance under ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})$ as:
$\mathcal{\rm cov}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})=\begin{pmatrix}\mathcal{C}^{vv}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})&\mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})\\ \mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})^{\top}&\mathcal{C}^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}})\end{pmatrix}.$
We use similar notation (${\rm mean}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}}$) for the mean under ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}.$ The ensemble Kalman filter from [evensen1994sequential] then has the form, for $n\in\mathbb{Z}^{+}$,
@@
$v_{n+1}^{(j)}$ $=\widehat{v}_{n+1}^{(j)}+\mathcal{C}^{vy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1}\bigr{)}\mathcal{C}^{yy}\bigl{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1}\bigr{)}^{-1}\bigl{(}{y}^{\dagger}_{n+1}-\widehat{y}_{n+1}^{(j)}\bigr{)},$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1}$ $=\frac{1}{J}\sum_{j=1}^{J}\delta_{\bigl{(}\widehat{v}_{n+1}^{(j)},\widehat{y}_{n+1}^{(j)}\bigr{)}}.$
@@
Note that the components of ${\rm cov}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ involve empirical learning of the covariances $\Sigma$ and $\Gamma$ of the $\{\xi_{n}^{(j)}\}_{j=1}^{J}$ and $\{\eta_{n+1}^{(j)}\}_{j=1}^{J}$ respectively. It is possible, and indeed often desirable, to directly input the matrices $\Sigma$ and $\Gamma$, only using empirical estimation for the covariance based on the $\{\Psi(v_{n}^{(j)})\}_{j=1}^{J}$ and $\{h(\widehat{v}_{n+1}^{(j)})\}_{j=1}^{J}.$ In particular to use
$\mathcal{C}^{vy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ $\approx C^{vh,J},$ (4.1a)
$\mathcal{C}^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ $\approx C^{hh,J}+\Gamma$ (4.1b)
where $C^{vh,J}$ is the empirical cross-correlation between the $\{\Psi(v_{n}^{(j)})\}_{j=1}^{J}$ and the $\{h(\widehat{v}_{n+1}^{(j)})\}_{j=1}^{J}$, whilst $C^{hh,J}$ is the empirical correlation of the $\{h(\widehat{v}_{n+1}^{(j)})\}_{j=1}^{J}$. The actual definition of $C^{yy}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}^{{\rm EK},J}_{n+1})$ instead uses an empirical approximation $\Gamma_{n+1}$ of $\Gamma$ formed from $\{\eta_{n+1}^{(j)}\}_{j=1}^{J}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{{\rm EK},J}_{n}=\frac{1}{J}\sum_{j=1}^{J}\delta_{v_{n}^{(j)}},$
noting that for $n=0$ we choose $v_{0}^{(j)}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{0}$ i.i.d. . Practitioners like this methodology because it assigns equal weights to the particles and cannot suffer from the weight collapse arising in the particle filter. Furthermore, in the setting where $\Psi,h$ are both linear, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ is Gaussian, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}^{{\rm EK},J}_{n}$ converges to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ as $J\to\infty$, at the Monte Carlo rate [le2009large, mandel2011convergence]. In this linear setting, with Gaussian noise and initial condition, the problem can be solved explicitly by the Kalman filter [kalman1960new]; this is a nonautonomous and nonlinear dynamical system for the mean and covariance of the (in this setting) Gaussian ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}.$ However forming and propagating $d\times d$ covariances, when $d\gg 1$, is impractical; ensemble methods, on the other hand, operate by computing a $J\times d$ low-rank approximation of the covariance and can be used when $d\gg 1.$ Thus particle methods are of some practical value even in this linear Gaussian setting, avoiding the need to work with large covariances when $d\gg 1.$